OpenStudy (anonymous):

The radius of a puddle is claimed to be 12.0 inches, correct to within 0.01 inch. Use linear approximation to estimate the resulting error, measured in square inches, in the area of the puddle. A) plus or minus 0.24pi B) Plus or minus 0.024pi C) plus or minus 0.06pi D)Plus or minus 0.0004166

OpenStudy (anonymous):

@ganeshie8 can u plz help ? :)

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@sammixboo @pooja195 can u plz help ?:)

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@Hero can u plz help? :)

OpenStudy (anonymous):

@CGGURUMANJUNATH can u plz help ? :)

OpenStudy (anonymous):

@freckles can u plz help? :)

OpenStudy (snowsurf):

You know the formula for any circular object is \[A=\pi r ^{2}\] Now what you do is take its derivative. \[dA=2\pi r dr\]where r is the radius and dr is the tiny difference. So plugging in values you are given \[dA = 2 \pi (12)(0.01)\] Use a calculator and you will get 0.24 pi.

OpenStudy (freckles):

http://www.zweigmedia.com/RealWorld/calctopic1/linearapprox.html it looks like you can use: \[\Delta y= \approx \Delta x \cdot \frac{dA}{dr}_{x=12}\]

OpenStudy (freckles):

r=12*

OpenStudy (freckles):

are actuallly snow has all the work above

OpenStudy (anonymous):

thank u all! 2 best answers :)